The rule of 72 is a rule of thumb for finding how long it takes money at interest to double:� If r is the annual interest rate, then the doubling time is approximately 72/(100r) years.

(a) Calculate the balance at the end of the predicted doubling time for each $1000, with annual compounding, for the small growth rates of 3%, 4%, and 6%
b) Repeat part (a) for the intermediate interest rates of 8% and 9%.

(c) Repeat part (a) for the larger interest rates of 12%, 24%, and 36%.

(d)What do you conclude about the rule of 72?

Answer by Theo(13342)

f = p * (1+r)^n

f = future value
p = present value
r = interest rate per time period
n = number of time periods.

when you compound annually, r is the annual interest rate and n is the number of years.

if you divide both sides of the equation by p, you get:

f/p = (1+r)^n

if your present value doubles, then f/p = 2

you wind up with 2 = (1+r)^n

take the natural log of both sides of this equation to get:

ln(2) = ln(1+r)^n

since log(a^b) = b*log(a), your equation becomes:

ln(2) = n*ln(1+r).

solve for n to get n = ln(2) / ln(1+r).

when r = .03, this becomes n = ln(2) / ln(1.03)
when r = .04, this becomes n = ln(2) / ln(1.04)
when r = .06, this becomes n = ln(2) / ln(1.06)

solve for n and you get:

when r = .03, n = 23.45
when r = .04, n = 17.67
when r = .06, n = 11.90

using the rule of thumb, .....

when r = .03, n = 72/3 = 24
when r = .04, n = 72/4 = 18
when r = .06, n = 72/6 = 12

that's a pretty close estimate to the actual calculations.

rule of 72 looks like it works pretty good.